# The Double-Ellipsoid Geometry of CLIP

### v1

@@ 6 False Negatives and Conformity
We show here how the geometric structure of the embedding, found in previous sections, is advantageous when false negatives exist. We then propose the notion of conformity, which appears to play a major role in forming the distribution of the latent space.
In contrastive learning, a well known problem is the phenomenon of false negatives. Essentially, concepts which are not dedicated pairs have similar meaning. Thus, they do not represent well negative examples and should not be embedded far apart. This problem appears for both single and multi-modality cases. There have been several suggestions to solve this problem by introducing new learning procedures or proposing alternative contrastive losses [4, 8].
In the case of CLIP, however, the training is performed with a contrastive loss which does not take into consideration directly the false negative problem. We claim that the issue is mitigated, to some extent, by a specialized embedding. In classification and segmentation, as uncertainty grows, the network yields softer values, representing the lower membership probability to the class. In Fig. 7 (top) a typical segmentation score of semantic segmentation is shown, where the boundary of the sheep is blurred, due to reduced confidence. We refer to this as semantic blur. For a contrastive network, when false negatives are suspected we expect lower confidence and a more blurred response. In the case of a high dimensional sphere centered at the origin, blurring is not trivial. Small perturbations yield large disparities in cosine distance. We show below how a sphere not centered at the origin can mitigate this.
Figure 7: Top: blur of the segmentation score (right), typical in sematic segmentation, as the uncertainty of object-membership grows. Bottom: the similarity histograms of normally distributed samples are shown for the mean vector (blue) and a vector furthest from the mean (“extreme”, orange). The plots are done for a sphere centered near the origin (left) and for one with a center far from the origin (right). See full details of the experiment in the text. In the case of contrastive learning, blur can be achieved, as uncertainty grows, by having a sphere centered far from the origin. The desired blur can be controlled by positioning the embedded vector closer to the sphere center (blur) or farther from it (sharp).
[…]
@@
Figure 8: High and low conformity of natural images from MS-COCO [36]. Low-conformity images often depict unique, distinguishable individuals or objects, whereas high-conformity images capture common scenes that could be found anywhere.
@@ 6.1 Conformity
In order to validate our prediction, we first formalize the term common themes, by defining a new notion, termed conformity.
@@
To provide more intuition, we present examples of high and low conformity from MS-COCO in Fig. 8, as well as on ImageNet-a [21] and ImageNet-R [20] in the Supp. Following our prediction above, we propose a surrogate measure of conformity (which is much faster to compute). The estimation uses the following definition.
@@
where $a$ and $b$ are scalars determined by the specific embedding.
Figure 9: Conformity. Estimated conformity $\hat{C}$, Eq. 11, against conformity $C$, Eq. 10, on MS-COCO [36]. The correlation is almost perfect. We can thus use the proposed estimated conformity reliably to quantify how common a sample is. Employing $\hat{C}$ is much faster, using a single mean vector $m$, without the need to store an entire set of representative image-text pairs. We also show a few caption examples of high and low conformity, with the same number of words (8). More exotic captions have lower conformity.
@@ 6.2 Modality gap assists in distribution matching
We now aim to provide a reason that can justify the presence of the well known modality gap [35]. Our rationale for that phenomenon is as follows.
The same incentive of having a mean not centered at the origin applies for both image and text modalities. However, in a single image-pair instance the uncertainty for each modality may differ (see Fig. 10). The same arguments as before promote uncertain instances to be near the mean and certain ones to be far from it. If both image and text of a pair are embedded at the same location - we may get contradicting requirements. Having separate embeddings for text and image allows to control the uncertainty of each instance for each modality. More generally, we would like to match the distribution of the conformity of both modalities. In Fig. 11 we show the KL-divergence of the conformity distribution as a function of $\alpha$, a parameter controlling the distance of the mean from the origin, as in Eq. 9, see illustration in Fig. 6. We show that the best distribution match is near $\alpha=0$, that is, in the current embedding of CLIP.
Figure 11: Matching the distribution of conformity for image and text by introducing a modality gap. The parameter $\alpha$ controls the shift from the origin for images (as in Fig. 6). Both image and text conformity distributions are aligned when $\alpha\approx 0$, i.e. without any shift from the trained setting. $KL_{\alpha=0}\approx 0.14$ yields a good distribution matching.
Figure 12: Vertical SLERP (vSLERP) - Optimization-free, spontaneous semantic editing. Images interpolated with vSLERP preserve the same object with slight pose variations, while roughly maintaining backgrounds. The interpolation magnitude is controlled by $\alpha$.
Figure 13: Conformity analysis of Captioning and Image Synthesis. Image Synthesis (top): Glide [41] tends to generate more common images, often struggling with fine-grained details, whereas unCLIP [45] generates images with more delicate details, aligning more closely with the distribution of natural images. Though there is still room for improvement. Captioning (bottom): ClipCap [40] generates captions with higher conformity (i.e., more common descriptions) compared to human-annotated captions. In contrast, Caption Reward [7] produces captions with lower conformity, even surpassing human annotations in terms of uniqueness.
@@ 7.1 Conformity as a measure of expressiveness
We propose to use (estimated) conformity as a metric to assess the richness of generative methods. Specifically, we measure the conformity of images generated from MS-COCO captions by unCLIP [45] and Glide [41], visualizing the results in Fig. 13. Glide-generated images exhibit high conformity, indicating less detail and diversity, whereas unCLIP images are more varied and detailed. However, both models fall short of matching the diversity found in real images.
Similarly, we evaluate captioning methods by measuring conformity in captions generated using ClipCap [40] and Caption Reward [7]. ClipCap tends to produce common and trivial captions, while Caption Reward generates captions with diversity surpassing even human annotations.
@@ 7.2 Unguided, training-free semantic generation
The unCLIP framework [45] introduces an image interpolation technique using spherical linear interpolation (SLERP) to transform a source image into a target image gradually. While this method produces visually appealing results, it often fails to preserve the same instance throughout interpolation, instead generating random instances.
In Fig. 12, we show images generated by an extension of SLERP, which we term as vertical SLERP (vSLERP), defined as:
@@
For brevity, $m_{i}$ and $v_{i}$ are referred to as $m$ and $v$. With a fixed $\Theta=\Theta_{0}$, adjusting $\alpha$ allows controlled manipulation of the same instance (values in the range $-2\leq\alpha\leq 2$ tend to produce visually appealing images). This approach parallels real-image editing techniques; however, unlike methods relying on text inversion [16, 13, 39] or test-time optimization [27], which are computationally heavy, vSLERP requires no training or optimization, thus, highly efficient.
@@ 8 Discussion and Conclusion
This paper analyzes the raw CLIP embeddings before projection onto a unit hypersphere, revealing that each modality’s distribution forms a distinct, shifted ellipsoid. These ellipsoids differ in properties such as their center of mass and radii and are disjoint, hence the modality gap phenomenon [35]. Upon projection onto a unit hypersphere, they occupy only a limited angular space, which appears as a narrow cone effect [35, 47]. We introduce the concept of conformity, a measure of similarity with the entire set, to gain further insights into the double-ellipsoid geometry. We observe that each modality has a unique conformity distribution, with optimal matching achieved when the ellipsoids are positioned away from the center. Using this framework, we evaluate the diversity of captioning and image synthesis methods and propose a training-free interpolation technique, vertical SLERP (vSLERP), for specific object interpolation.

### v2

@@ 6 False negatives and conformity
We demonstrate how the embedding geometry discussed earlier provides advantages in handling false negatives. Additionally, we introduce the concept of conformity, which plays a major role in forming the latent space distribution. A well-known challenge in contrastive learning is the presence of false negatives—pairs with similar meanings that are not dedicated pairs. Such samples should not be embedded far apart, as they fail to represent true negatives effectively. This issue arises in both single- and multi-modality settings and has been addressed by proposing new training procedures or alternative contrastive losses (Byun et al., 2024; Chuang et al., 2020). In CLIP, training uses a contrastive loss that does not explicitly address false negatives. However, we argue that this issue is partially mitigated by the embedding geometry. In classification and segmentation tasks, uncertainty typically results in softer predictions that reflect lower class membership probabilities. For example, Fig. 7 (top) illustrates a segmentation score where reduced confidence blurs the sheep’s boundary, a phenomenon we term semantic blur.
Figure 8: High and low conformity of MS-COCO. Low-conformity images often depict unique, distinguishable individuals or objects, whereas high-conformity images capture common scenes that could be found anywhere.
For contrastive networks, when false negatives are present, we expect lower confidence and a blurred response. On a high-dimensional sphere centered at the origin, such blurring is challenging, as small perturbations lead to large changes in cosine distance. We show that shifting the sphere away from the origin can effectively mitigate this issue. Concurrently, and closely related, Schrodi et al. (2024) discuss the relationship between entropy and the modality gap.
Blur through a non-origin centered sphere. Fig. 7 (bottom) illustrates the difference between origin-centered and non-origin-centered spheres through an experiment. We draw 1000 random vectors $v^{j}\in\mathbb{R}^{512}$, where each element follows an independent Gaussian distribution with unit standard deviation. In the first experiment (Fig. 7, bottom left), the sphere is centered at the origin with an empirical mean $m$ close to zero. The blue histogram shows $\cos(m,v^{j})$ for $j=1,\dots,1000$. We then identify the furthest vector from $m$, $v^{\text{far}}=\arg\min\cos(m,v^{j})$, and plot the histogram of $\cos(v^{\text{far}},v^{j})$ (orange), excluding $v^{\text{far}}$. In the second experiment (Fig. 7, bottom right), the sphere is centered at $(10,5,5,0,0,\dots)$, modeling three dominant features with a mean distinctly far from zero. The same trial is repeated. The results highlight a significant difference: for an origin-centered sphere, the distributions of cosine similarity for the mean and the extreme vector are similar. In contrast, for a non-origin-centered sphere, the mean vector exhibits much higher average similarity. This allows the network to embed vectors with uncertainty closer to the mean, enabling semantic blur—reduced contrast in the response. This analysis, supporting a non-zero mean, leads to the following prediction:
@@
Figure 9: Conformity. Estimated conformity $\hat{C}$, Eq. 11, against conformity $C$, Eq. 10, on MS-COCO (Lin et al., 2014). The correlation is almost perfect. We can thus use the proposed estimated conformity reliably to quantify how common a sample is. More exotic captions have lower conformity (all examples are of eight words).
Figure 11: Modality Gap matches conformity distributions. The parameter $\alpha$ controls the embedding offset from the origin (as shown in Fig. 6). When $\alpha\approx 0$, i.e., the trained setting, image and text conformity distributions align well, with $KL_{\alpha=0}\approx 0.14$ indicating good distribution matching.
@@ 6.1 Conformity
To validate Prediction 1, we first formalize the term common themes, by defining a new notion, termed conformity.
@@
To provide more intuition, we present examples of high and low conformity from MS-COCO in Fig. 8, as well as on ImageNet-a (Hendrycks et al., 2021b) and ImageNet-R (Hendrycks et al., 2021a) in the Appendix. Following our prediction above, we propose a surrogate measure of conformity (which is much faster to compute). The estimation uses the following definition.
@@
where $a$ and $b$ are scalars determined by the embedding.
@@ 6.2 Modality gap assists in distribution matching
We now aim to provide a reason that can justify the presence of the well known modality gap (Liang et al., 2022). Our rationale for that phenomenon is as follows. The same incentive of having a mean not centered at the origin applies for both image and text modalities. However, in a single image-pair instance the uncertainty for each modality may differ (see Fig. 10). The same arguments as before promote uncertain instances to be near the mean and certain ones to be far from it. If both image and text of a pair are embedded at the same location - we may get contradicting requirements. Having separate embeddings for text and image allows to control the uncertainty of each instance for each modality. More generally, we would like to match the distribution of the conformity of both modalities. In Fig. 11 we show the KL-divergence of the conformity distribution as a function of $\alpha$, a parameter controlling the distance of the mean from the origin, as in Eq. 9, see illustration in Fig. 6. We show that the best distribution match is near $\alpha=0$, i.e., in the current embedding of CLIP.
Figure 12: Vertical SLERP (vSLERP) enables optimization-free, semantic editing. Interpolated images preserve the object with pose variations and roughly maintain backgrounds, with interpolation magnitude controlled by $\alpha$.
Figure 13: Conformity analysis of captioning and image synthesis. Image Synthesis (top): Glide generates more common images with less fine detail, while unCLIP creates more detailed images closer to natural distributions. Captioning (bottom): ClipCap produces more common captions, while Caption Reward generates more unique captions, even surpassing human annotations.
@@ 7.1 Conformity as a measure of expressiveness
We propose using conformity as a metric to assess generative method diversity. We measure conformity in images generated from MS-COCO captions by unCLIP (Ramesh et al., 2022) and Glide (Nichol et al., 2021), as shown in Fig. 13. Glide-generated images exhibit high conformity, indicating low detail and diversity, while unCLIP images are more varied and detailed. Both models, however, lack the diversity seen in real images. Similarly, we evaluate captioning methods by measuring conformity in captions generated by ClipCap (Mokady et al., 2021) and Caption Reward (Cho et al., 2022). ClipCap produces common captions, while Caption Reward generates diverse captions that even surpass human annotations.
@@ 7.2 Unguided, training-free semantic generation
The unCLIP framework (Ramesh et al., 2022) introduces an image interpolation technique using spherical linear interpolation (SLERP) to transform a source image into a target image gradually. While this method produces visually appealing results, it often fails to preserve the same instance along interpolation, instead generating random instances.
In Fig. 12, we show images generated by an extension of SLERP, which we term as vertical SLERP (vSLERP):
@@
For brevity, $m_{i}$ and $v_{i}$ are referred to as $m$ and $v$. With a fixed $\Theta=\Theta_{0}$, adjusting $\alpha$ allows controlled manipulation of the same instance. This approach parallels real-image editing techniques; however, unlike methods relying on text inversion (Han et al., 2024; Gal et al., 2022; Mokady et al., 2023) or test-time optimization (Kawar et al., 2023), which are computationally heavy, vSLERP requires no training or optimization, thus, highly efficient.
@@ 8 Discussion and Conclusion
The paper examines the primary CLIP embedding, prior to projection onto the unit sphere, revealing that each modality forms a distinct, shifted ellipsoid with unique centers and radii. This geometry is the source of the modality gap and narrow cone phenomena (Liang et al., 2022; Schrodi et al., 2024; Afham et al., 2022), previously observed on the unit sphere embedding. We introduced conformity, a measure of similarity of an instance with an entire representative set. Our analysis shows that each modality exhibits a unique conformity distribution, with optimal alignment achieved when the ellipsoids are shifted from the origin. This provides a useful tool for assessing the diversity of captioning and image synthesis methods. Finally, we propose vertical SLERP (vSLERP), a training-free interpolation technique for specific object interpolation.
Acknowledgements
We would like to acknowledge support by the Israel Science Foundation (Grant 1472/23) and by the Ministry of Science and Technology (Grant No. 5074/22).
Impact Statement
Our work advances machine learning by improving the geometrical understanding of CLIP’s latent space. The findings may influence downstream tasks, though specific societal consequences do not need further emphasis here.
Appendix A Enlraged Visualizations
In Figure 15 and Figure 16, we provide the same visualizations as in the main paper, but enlraged, to enhance visibility. CLIP of higher dimension. We also show some results for CLIP with ViT-L/14 encoders, $n=768$. In Figure 17 we show the distinct different statistics of image and text, mostly appearing in several pronounced features. Here as well, linear separation (100% classification accuracy) can be reached with only two features. In Figure 18 we show that the embedding can also be modeled as two separate thin shell ellipsoids for image and text.
Appendix B Statistical Analysis
We provide here the definitions of log concave distributions and isotropic random vectors, notions which are used in Section 4 of the main paper.
Definition 3 (Log concave distribution).
A log concave distribution in $R^{n}$ has a density $p$ which admits, $\forall x,y\in R^{n},\lambda\in[0,1]$,
$p(\lambda x+(1-\lambda)y)\geq p(x)^{\lambda}p(y)^{1-\lambda}.$
The above definition is equivalent to stating that the logarithm of the density function is concave $\log p(\lambda x+(1-\lambda)y)\geq\lambda\log p(x)+(1-\lambda)\log p(y)$. Many well-known distributions admit this property, such as normal and multivariate normal distributions, exponential, Laplace, chi, Dirichlet, gamma and more.
Definition 4 (Isotropic random vector).
A random vector $x\in R^{n}$ is isotropic if $\operatorname{\mathbb{E}}[x]=0$ and $\Sigma=I$, where $\Sigma$ is the covariance matrix of $x$ and $I$ is the identity matrix.
Figure 14: Norm distribution. While norm magnitudes are disregarded during training due to the normalization inherent in cosine similarity, they still capture meaningful semantic information.
We give below additional analysis related to applying a linear transformation that turns each ellipsoid into a sphere. This process is termed sphering or whitening. For lack of space, this part did not get into the main paper. However, we believe this analysis is of sufficient merit to be presented here.
Figure 15: Enlarged plots from Section 4.
Figure 16: Enlarged plots from Section 4.
Figure 17: Enlarged plots for CLIP embedding of $n=768$. There are dominant features with clearly different distribution between image and text. Both modalities can be separated (with perfect accuracy) by a linear SVM classifier based on only 2 features. With respect to separability (bottom), there are 20 features with value above 1.
Figure 18: CLIP $n=768$, thin shell phenomenon. We can observe similar geometry (as in the case of $n=512$) of two tilted ellipsoids, one for each modality, not centered at the origin.
Appendix C Additional Experiments and Visualizations
C.1 Conformity
High- and Low-Conformity Images. We provide additional visualizations of high- and low-conformity images across various datasets. Figure 19 illustrates examples of sketches from ImageNet-R, while Figure 20 showcases examples from ImageNet-A. Both datasets contain out-of-distribution examples: ImageNet-A emphasizes natural adversarial images, while ImageNet-R features renditions of objects, such as origami or sketches.
From these visualizations, we observe that high-conformity images tend to contain less information. Sketches are simpler, and natural images often feature large uniform backgrounds or repetitive structures. In contrast, low-conformity images frequently include substantial text, while natural images exhibit collages of objects with unique or diverse colors.
C.2 Reaffirming loss and conformity matching experiments
We revisit the loss experiment presented in Fig. 6 of the main paper and the conformity matching experiment shown in Fig. 11. To further validate our findings, we conduct these experiments under two alternative settings.
First, we shift the text ellipsoid instead of the image ellipsoid, applying the following transformation:
$v_{t}^{j^{\prime}}=v_{t}^{j}-\alpha\cdot m_{t}\quad\forall j\in M,$ (13)
where the values of $v_{i}$ remain unchanged. The results of this experiment are presented in Figure 21.
In the second setting, we align both the image and text ellipsoids at the origin by applying the following transformations:
$v_{t}^{j^{\prime}}=v_{t}^{j}-\alpha\cdot m_{t},\quad v_{i}^{j^{\prime}}=v_{i}^{j}-\alpha\cdot m_{i}\quad\forall j\in M.$ (14)
Here, for $\alpha=0$, the ellipsoids remain in their optimal positions after training, while for $\alpha=1$, both ellipsoids are shifted to the origin as in Figure 22.
Both experiments reaffirm that the current positioning of the ellipsoids yields optimal results in terms of loss and conformity matching. These findings further support our claims across different alignment scenarios.
C.3 vSLERP
Here, we provide additional examples of vSLERP, shown in Figure 23 and Figure 24. As discussed in the main paper, the standard SLERP process typically generates interpolated images representing different objects or individuals. In contrast, our proposed vSLERP method produces diverse variations of the same object.
Figure 19: High and low conformity of sketches from ImageNet-R. Images with high conformity tend to be simpler and cleaner, while low-conformity images often feature complex details covered by large portions of text descriptions.
Figure 20: Conformity on ImageNet-a. It is possible that high conformity images are with more unique colors, perhaps contains people or text, whereas low conformity images tends to contain low amount of information.
Figure 21: Shifting text ellipsoid only. Conformity distribution matching and loss experiments when shifting text ellipsoid only as in Equation 13
Figure 22: Shifting both ellipsoids. Conformity distribution matching and loss experiments when shifting both text and image ellipsoids as in Equation 14.
Figure 23: vSLERP lamp to vase.
Figure 24: vSLERP Kevin Durant to Lebron James.